Circular No. 7498

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 S2 (LINEAR) F. Shelly and R. Huber, MIT Lincoln Laboratory, report the discovery

by LINEAR of a comet with a tail in p.a. ∼ 280^{o}, as given below. Confirmation of cometary appearance on Sept. 29.1 UT has been received from P. Pravec and P. Kušnirák at Ondřejov (coma diameter 0′.2 with a 0′.2 tail in p.a. 260^{o}) and from J. Tichá and M. Tichý at Kleť (object diffuse with 14′′ tail in p.a. 265^{o}).

2000 UT  α_{2000}  δ_{2000}  m_{1}
Sept.27.44573  7^{h}11^{m}15.^{s}46  +27^{o}43′48′′.5  18.6

Available astrometry and preliminary parabolic orbital elements (T = 2000 July 27.6 TT, q = 1.305 AU, i = 20^{o}.2) are published on MPEC 2000-S67.

SUPERNOVA 2000dn IN IC 1468 C. Yu and W. D. Li, University of California at Berkeley, on behalf of

the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Sept. 27.2 (mag ∼ 17.9) and 28.2 UT (mag ∼ 17.7), and located at α = 23^{h}05^{m}05^{s}.86, δ = −3^{o}12′01′′.0 (equinox 2000.0), which is 25′′.8 west and 15′′.2 north of the nucleus of IC 1468. A KAIT image of the same field taken on Sept. 16.2 showed nothing at the position of SN 2000dn (limiting mag ∼ 18.5).

X-RAY TRANSIENT IN M32 M. R. Garcia, S. S. Murray, F. A. Primini, and J. E. McClintock,

Smithsonian Astrophysical Observatory; and P. J. Callanan, University College, Cork, communicate: “We have discovered a bright x-ray transient apparently within M32, ∼ 40′′ southwest of the nucleus of M32. The transient is located at α = 0^{h}42^{m}40^{s}.7, δ = +40^{o}51′17′′.7 (equinox 2000.0; 90-percent error radius 2′′.0). The source luminosity is ∼ 10^{38} erg/s in the Chandra bandpass (0.2–10.0 keV), as seen in HRC and ACIS images from Sept. 11 and 21, respectively. A Chandra observation on Aug. 18 sets a pre-outburst upper limit on the source luminosity of 10^{36} erg/s. The next Chandra observations of this field will occur during the weeks starting Oct. 9 and 16. The scaling of x-ray transients within our own galaxy (e.g., GRO J0422+32 and A0620–00) to the distance of M32 indicates that an optical counterpart may be as bright as U ∼ 24.”

2000 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green